OpenStudy (anonymous):

Evaluate y/x x=5 y= -4 m/n m= -2 n=8 2m-11/n m=7 n=14

ganeshie8 (ganeshie8):

simply plugin the given values in the given expression, and simplify if possible

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

@love10129151 @tester97 @Haseeb96 help ?

OpenStudy (haseeb96):

just put the values in that equation i am going to give u an example x=9, y=3 2x/y now just put these above a value in that equation 2(9)/3 = 18/3 =6 is the correct answer like this solve your questions